# Design Stability in Adaptive Experiments: Implications for Treatment Effect Estimation

### v1

@@ 4 Main Results
$p_{i}\xrightarrow{p}p^{\star}.$ (10)
Although the notion of strong design stability is intuitive, it is not satisfied by several popular designs. A concrete example is Efron’s biased coin design [3], which enforces balance between treatment and control assignments. Fortunately, Definition 2 can be relaxed so that, even if a design is not stable in the strong sense, central limit theorems for the IPW and AIPW estimators may still hold under weaker regularity conditions. This motivates the following weaker notion of stability.
@@
$\frac{1}{N}\sum_{i=1}^{N}\frac{1}{p_{i}}\xrightarrow{p}\frac{1}{p_{1}^{\star}}\qquad\text{and}\qquad\frac{1}{N}\sum_{i=1}^{N}\frac{1}{1-p_{i}}\xrightarrow{p}\frac{1}{1-p_{2}^{\star}}.$ (11)
The tradeoff between these two stability notions becomes apparent when estimating the asymptotic variance of our estimators to construct confidence intervals for the ATE. While variance estimators can be constructed in a completely data-dependent manner under strong stability, weak stability requires additional restrictions (see Theorems 3 and 6 for more details).
@@ 4.1 The IPW estimator
The proof of Theorem 1 proceeds by rewriting the centered and scaled IPW estimator as a sum of martingale difference terms. This representation allows us to apply the martingale central limit theorem [32, Chapter 3]. We establish unbiasedness by verifying that each summand has zero conditional mean, and then compute the conditional variance, which converges under both stable and weakly stable designs to the stated asymptotic variance. Finally, we check the Lindeberg condition, ensuring that the contribution of large deviations vanishes. Together, these steps yield asymptotic normality of the IPW estimator with the asymptotic variance given in (13). Refer to Supplementary material 7.1 for detailed proof.
Having established the asymptotic normality of the IPW estimator, we next construct confidence intervals for the average treatment effect $\bar{\tau}$. This, in turn, requires estimation of the asymptotic variance $\mathrm{V}^{\mathrm{IPW}}$. First, we consider that the design is strongly stable in the sense of Definition 2, and assume that $p^{*}$ is known (which is the case in our illustrative example on strongly stable designs). To estimate $\mathrm{V}^{\text{IPW}}_{\text{strong}}$ in (13), we must estimate $m^{2}_{0}$, $m^{2}_{1}$, and $m_{01}$. While obtaining consistent estimators of $m^{2}_{0}$ and $m^{2}_{1}$ under strong stability is straightforward, the cross-moment term $m_{01}$ cannot be estimated from the observed outcomes without additional assumptions, as only one potential outcome is observed for each unit. To address this problem, we apply the Cauchy–Schwarz inequality to obtain $|m_{01}|\leq m_{0}m_{1}$, leading to the conservative variance estimator:
@@
where $N_{1}=\sum_{i=1}^{N}K_{i}$ and $N_{0}=N-N_{1}$.
The variance estimator in (14), which incorporates the estimators $\widehat{m}_{0}^{2}$ and $\widehat{m}_{1}^{2}$ defined in (15), is consistent when the potential outcomes are additive on a log-scale, that is, satisfy (4) and has an asymptotic positive bias otherwise. Hence, we obtain the following theorem.
@@
See Supplementary material 7.2 for a proof of the theorem.
@@
may be substituted for $p^{*}$ into the variance estimator $\widehat{V^{\text{IPW}}_{\text{strong}}}$ (14).
@@
$\frac{1}{N}\sum_{i=1}^{N}p_{i}\xrightarrow{p}\widetilde{p}$, (18)
@@
As in the strongly stable case, the variance estimator in (17), which incorporates the estimators $\widetilde{m}_{0}^{2}$ and $\widetilde{m}_{1}^{2}$ defined in (19) and (20), respectively, is consistent when the potential outcomes are additive on the log-scale; that is, when they satisfy (4). Thus, we obtain the following theorem.
@@
Refer to Supplementary material 7.3 for a proof of the theorem.
@@
Substituting the above estimators into (17), a conservative estimator of $\mathrm{V}^{\mathrm{IPW}}$ under weak stability when $p_{1}^{\star},p_{2}^{\star}$ and $\widetilde{p}$ are unknown is:
$\widehat{V^{\text{IPW}}_{\text{weak}}}=\widetilde{m}_{0}^{2}\left(\frac{1}{N}\sum_{i=1}^{N}\frac{1}{1-p_{i}}-1\right)+\widetilde{m}_{1}^{2}\left(\frac{1}{N}\sum_{i=1}^{N}\frac{1}{p_{i}}-1\right)+2\widetilde{m}_{0}\widetilde{m}_{1},$ (22)
@@
Since we have constructed conservative variance estimators of $\mathrm{V}^{\mathrm{IPW}}$, for any target level $\alpha\in(0,1)$, asymptotically conservative confidence interval for $\bar{\tau}$, that is, one with coverage at least $(1-\alpha)$ can be constructed,
$\lim_{N\rightarrow\infty}\mathbb{P}\left(\bar{\tau}\in\left[\widehat{\tau}_{\textsc{IPW}}-z_{1-\alpha/2}\cdot\sqrt{\frac{\widehat{V}}{N}},\;\;\widehat{\tau}_{\textsc{IPW}}+z_{1-\alpha/2}\cdot\sqrt{\frac{\widehat{V}}{N}}\right]\right)\geq 1-\alpha,$
where $\widehat{V}=\widehat{V^{\text{IPW}}_{\text{strong}}}$ or $\widehat{V^{\text{IPW}}_{\text{weak}}}$, and $z_{1-\alpha/2}$ denotes the $(1-\alpha/2)$th quantile of the standard normal distribution. Moreover, if the potential outcomes satisfy the log-additive treatment-effect model (4), the inequality holds with equality, yielding asymptotically exact coverage.
@@ 4.2 The AIPW-type estimator
The proof of Theorem 4 requires a different strategy from that of Theorem 1, since the AIPW estimator is not directly amenable to martingale central limit theorem. To handle this, we introduce a proxy estimator that is analytically more tractable and can be expressed as sum of a martingale difference sequence, allowing the martingale central limit theorem to establish its asymptotic normality. The key step is then to show that the difference between the proxy and the actual AIPW estimator is asymptotically negligible, using variance bounds and Hájek’s lemma (see Supplementary material 6). This ensures that the asymptotic distribution of the AIPW estimator coincides with that of the proxy, yielding the stated central limit theorem with variance given in Theorem 4. See Supplementary material 7.4 for detailed proof of the above theorem.
@@
Note that $\mathrm{V}^{\mathrm{AIPW}}\leq\mathrm{V}^{\mathrm{IPW}}$; that is, $\widehat{\tau}_{\textsc{AIPW}}$ is more efficient than $\widehat{\tau}_{\textsc{IPW}}$. This fact clearly establishes the superiority of the AIPW estimator over the IPW estimator in finite population design-based inference under the adaptive assignment mechanism defined in (7).
@@
A proof of this theorem is given in Supplementary material 7.5.
@@
As discussed previously, the cross-moment term $\sigma_{01}$ cannot be estimated without additional assumptions, since it depends on both potential outcomes for all units. We therefore bound it from above using the Cauchy–Schwarz inequality. Consequently, asymptotic variance $\mathrm{V}^{\text{AIPW}}_{\text{weak}}$ can be conservatively estimated by:
@@
As in the strong stability case, this estimator is consistent when the potential outcomes satisfy generalized treatment effect additivity according to Definition 1. The above discussion is summarized in the following theorem.
@@
See Supplementary material 7.6 for the proof.
@@
As in Section 4.1, for any target level $\alpha\in(0,1)$, asymptotically conservative confidence intervals for $\bar{\tau}$ may be constructed around $\widehat{\tau}_{\textsc{AIPW}}$ using the variance estimators (26) and (30) for strongly stable and weakly stable designs, respectively. If the limiting values of the probabilities are unknown, their counterparts suggested in Remark 6 may be used. All of these intervals are asymptotically conservative, but attain exact asymptotic coverage when the potential outcomes satisfy generalized treatment effect additivity (3).
@@ 5 Some illustrative applications
In this section, we illustrate Theorems 1-6 through two adaptive designs: a strongly stable design, Wei’s adaptive coin design [2], and a weakly stable design, Efron’s biased coin design [3]. We further complement the theoretical results with numerical simulations that demonstrate the validity of our approach.
@@ 5.1 Wei’s Adaptive Coin Design
We begin with Wei’s adaptive coin design [2], which reduces relative imbalance between treatment and control allocations. Formally, let $m_{k}$ and $n_{k}$ denote, respectively, the numbers of treatment and control units among the first $k$ subjects. Define the treatment-control imbalance as $D_{k}=m_{k}-n_{k},$ and the corresponding normalized imbalance $R_{k}=\frac{D_{k}}{k},$ which measures the average difference between the treatment and control groups up to stage $k$.
Under Wei’s adaptive coin design, the $i$th subject is assigned to treatment with probability
@@
Intuitively, when the trial is in its early stages, the number of units in each group can differ substantially in relative terms; the design then shifts $p_{i}$ away from $\tfrac{1}{2}$ to favor the smaller group and reduce imbalance. As the sample size grows, any absolute difference in group sizes becomes small relative to the total number of units, causing $R_{i-1}$ to shrink and $p_{i}$ to converge to $\tfrac{1}{2}$. The following Lemma, which is a direct consequence of [2, Theorem 1], establishes strong stability of the truncated version of Wei’s design (33), making Theorems 1245 directly applicable.
@@
See Supplementary material 8.1 for the proof of Lemma 1.
Substituting $p^{\star}=\tfrac{1}{2}$ into the expressions for $\mathrm{V}^{\mathrm{IPW}}$ and $\mathrm{V}^{\mathrm{AIPW}}$ in Theorems 1 and 4 gives the limiting variances of the IPW and AIPW estimators, $\widehat{\tau}_{\textsc{IPW}}$ and $\widehat{\tau}_{\textsc{AIPW}}$, respectively, under Wei’s design:
@@
Since $p^{\star}$ is known and fixed at $\tfrac{1}{2}$, it can be directly plugged into the variance estimators $\widehat{V^{\text{IPW}}_{\text{strong}}}$ in (14) and $\widehat{V^{\text{AIPW}}_{\text{strong}}}$ in (26), yielding the following conservative estimators for the IPW and AIPW variances:
@@
where $\widehat{m}^{2}_{\ell}$ and $\widehat{\sigma}^{2}_{\ell}$, $\ell\in\{0,1\}$, are as defined in (15) and (27), respectively. We can now use these variance estimators in place of $\widehat{V^{\text{IPW}}_{\text{strong}}}$ and $\widehat{V^{\text{AIPW}}_{\text{strong}}}$ to construct conservative confidence intervals for $\bar{\tau}$. Recall that the interval based on $\widehat{V^{\text{IPW}}_{\text{Wei}}}$ attains exact asymptotic coverage when the potential outcomes satisfy additivity on the log scale (4), whereas the interval based on $\widehat{V^{\text{AIPW}}_{\text{Wei}}}$ attains exact asymptotic coverage under generalized treatment effect additivity, as defined in (3).
(a) IPW estimator under Non-Additive case
(b) AIPW estimator under Non-Additive case
(c) IPW estimator under Additive case
(d) AIPW estimator under Additive case
(e) IPW estimator under Log-Additive case
(f) AIPW estimator under Log-Additive case
@@
Next, we evaluate the performances of the IPW and AIPW estimators under Wei’s adaptive coin design through simulation studies. We consider three data-generating mechanisms: (a) a general, non-additive outcome model; (b) the additive model in equation (2); and (c) the log-additive model in equation (4).
In the non-additive setting, the potential outcomes $(Y_{i}(0),Y_{i}(1))$ are drawn from a bivariate normal distribution with mean vector $(0,1)^{\text{T}}$ and variance–covariance matrix
$\begin{bmatrix}1&0.3\\ 0.3&1\end{bmatrix}$,
with support restricted to [3,3][-3,3] to ensure bounded outcomes.
In the additive setting, the control potential outcomes are drawn from a normal distribution with mean 0 and variance 1, truncated to [3,3][-3,3], and the treatment outcomes are defined by $Y_{i}(1)=Y_{i}(0)+\tau$ with $\tau=10$.
In the log-additive setting, the control potential outcomes are drawn from a normal distribution with mean 10 and variance 1, truncated to $[7,13]$, and the treatment outcomes are defined by $Y_{i}(1)=c\,Y_{i}(0)$ with $c=2$.
Treatment assignments $\mathbf{K}$ are generated according to Wei’s sequential randomization scheme, with assignment probabilities $p_{i}=f(R_{i-1})=\tfrac{1-R_{i-1}}{2}$, where $R_{i-1}$ denotes the normalized treatment–control imbalance prior to assigning the $i$th unit, and the truncation parameter is set to $\delta=0.01$. Each simulation involves N=5, [^0] N=5{,}000 units and is repeated 20,00020{,}000 times. For each replication, confidence intervals are constructed using the proposed methodology, and empirical coverage is evaluated across 20 nominal levels ranging from 0.75 to 0.99.

### v2

@@ 4 Main Results
$p_{i}\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}p^{\star}.$ (10)
Although the notion of strong design stability is intuitive, it is not satisfied by several popular designs. A concrete example is Efron’s biased coin design [Efron1971-yh], which enforces balance between treatment and control assignments. Fortunately, Definition 2 can be relaxed so that, even if a design is not stable in the strong sense, central limit theorems for the IPW and AIPW estimators may still hold under weaker regularity conditions. This motivates the following notion of weak stability.
@@
$\frac{1}{N}\sum_{i=1}^{N}\frac{1}{p_{i}}\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}\frac{1}{p_{1}^{\star}}\qquad\text{and}\qquad\frac{1}{N}\sum_{i=1}^{N}\frac{1}{1-p_{i}}\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}\frac{1}{1-p_{2}^{\star}}.$ (11)
Remark 1.
Technically, the inclusion probabilities can be viewed as a triangular array $\{p_{i,N}:1\leq i\leq N,\;N\geq 1\}$, which induces a sequence of estimands $\{\bar{\tau}_{N}\}_{N\geq 1}$ and, a sequence of estimators $\{\widehat{\tau}_{N}\}_{N\geq 1}$. Analysis for a fixed $N$ is referred to as finite-sample, whereas statements concerning the limiting behavior as $N\to\infty$ are termed asymptotic. Although our framework is asymptotic, most of our results are derived from finite-sample arguments and are interpreted through this asymptotic lens. For notational simplicity, we suppress the dependence on $N$ whenever no ambiguity arises.
Suppose strong stability holds, Since, by assumption, the inclusion probabilities are bounded away from both 0 and $1$, Slutsky’s theorem implies $\frac{1}{p_{i}}\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}\frac{1}{p^{\star}}$ and $\frac{1}{1-p_{i}}\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}\frac{1}{1-p^{\star}}$ and hence, $\frac{1}{N}\sum_{i=1}^{N}\frac{1}{p_{i}}\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}\frac{1}{p^{\star}}$ and $\frac{1}{N}\sum_{i=1}^{N}\frac{1}{1-p_{i}}\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}\frac{1}{1-p^{\star}}$, which shows that strong stability implies weak stability with $p_{1}^{\star}=p_{2}^{\star}=p^{\star}$. The tradeoff between these two stability conditions becomes apparent when estimating the asymptotic variance of our estimators to construct confidence intervals for the ATE. While variance estimators can be constructed in a completely data-dependent manner under strong stability, weak stability requires additional restrictions (see Theorems 3 and 6 for more details).
@@ 4.1 The IPW estimator
The proof of Theorem 1 proceeds by rewriting the centered and scaled IPW estimator as a sum of martingale difference terms. This representation allows us to apply the martingale central limit theorem [Hall1980-ny, Chapter 3]. We establish unbiasedness by verifying that each summand has zero conditional mean, and then compute the conditional variance, which converges under both strongly stable and weakly stable designs to the stated asymptotic variance. Finally, we check the Lindeberg condition, ensuring that the contribution of large deviations vanishes. Together, these steps yield asymptotic normality of the IPW estimator with the asymptotic variance given in (13). Refer to Section S8.1 of the Supplementary Material for detailed proof.
Remark 3.
It is worth noting that our stability definitions provide two distinct mechanisms to ensure that the random conditional variance converges in probability to a deterministic limit. As seen in the proof of Theorem 1, this conditional variance is
$V_{N}=\frac{1}{N}\sum_{i=1}^{N}\frac{p_{i}}{1-p_{i}}Y_{i}(0)^{2}+\frac{1}{N}\sum_{i=1}^{N}\frac{1-p_{i}}{p_{i}}Y_{i}(1)^{2}+\frac{2}{N}\sum_{i=1}^{N}Y_{i}(0)Y_{i}(1).$
Strong stability requires $p_{i}$ to converge in probability to fixed limit $p^{\star}$. Intuitively, the design asymptotically “freezes” as every term in the sum $V_{N}$ becomes constant. On the other hand, weak stability permits $p_{i}$ to fluctuate but requires the sample means of $N^{-1}\sum_{i=1}^{N}p_{i}^{-1}$ and $N^{-1}\sum_{i=1}^{N}(1-p_{i})^{-1}$ to converge in probability to some fixed limits $1/p_{1}^{\star}$ and $1/p_{2}^{\star}$. Intuitively, this acts as a Law of Large Numbers for the variance process. While individual weights $p_{i}/(1-p_{i})$ and $(1-p_{i})/p_{i}$ remain volatile, their average converges settles to a fixed constant. This ensures that the “total precision” of the experiment becomes predictable, satisfying the martingale CLT conditional variance stabilization condition without requiring the assignment algorithm to commit to a specific allocation.
Having established the asymptotic normality of the IPW estimator, we next construct confidence intervals for the average treatment effect $\bar{\tau}$. This, in turn, requires estimation of the asymptotic variance $\mathrm{V}^{\mathrm{IPW}}$. First, we consider that the design is strongly stable in the sense of Definition 2, and assume that $p^{\star}$ is known (which is the case in our illustrative example on strongly stable designs). To estimate $\mathrm{V}^{\text{IPW}}_{\text{strong}}$ in (13), we must estimate $m^{2}_{0}$, $m^{2}_{1}$, and $m_{01}$. While obtaining consistent estimators of $m^{2}_{0}$ and $m^{2}_{1}$ under strong stability is straightforward, the cross-moment term $m_{01}$ cannot be estimated from the observed outcomes without additional assumptions, as only one potential outcome is observed for each unit. To address this problem, we apply the Cauchy–Schwarz inequality to obtain $|m_{01}|\leq m_{0}m_{1}$, leading to the conservative variance estimator:
@@
where $N_{1}=\sum_{i=1}^{N}K_{i}$ and $N_{0}=N-N_{1}$. The variance estimator in (14), which incorporates the estimators $\widehat{m}_{0}^{2}$ and $\widehat{m}_{1}^{2}$ defined in (15), is consistent when the potential outcomes are additive on a log-scale, that is, satisfy (4) and has an asymptotic positive bias otherwise. Hence, we obtain the following theorem.
@@
Refer to Section S8.2 of the Supplementary Material for a proof of the theorem.
@@
may be substituted for $p^{\star}$ into the variance estimator $\widehat{V^{\text{IPW}}_{\text{strong}}}$, as defined in (14).
Remark 5.
$\widehat{V^{\text{IPW}}_{\text{strong}}}$ closely resembles the conservative variance estimator $\widehat{V}^{\prime}$ for the variance of the difference-in-means estimator [Ding2023-ku, Homework Problem 4.5]. Under strong stability, the effective sample sizes satisfy $N_{1}\approx Np^{\star}$ and $N_{0}\approx N(1-p^{\star})$ asymptotically. Substituting these approximations into the variance estimator $\widehat{V}^{\prime}$ in [Ding2023-ku, Homework Problem 4.5] yields the same functional form as our conservative variance estimator.
@@
$\frac{1}{N}\sum_{i=1}^{N}p_{i}\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}\widetilde{p}$, (18)
@@
Some intuition and an example illustrating when weak stability holds but (18) does not, is provided in Section S13 of the supplementary material. As in the strongly stable case, the variance estimator in (17), which incorporates the estimators $\widetilde{m}_{0}^{2}$ and $\widetilde{m}_{1}^{2}$ defined in (19) and (20), respectively, is consistent when the potential outcomes are additive on the log-scale; that is, when they satisfy (4). Thus, we have the following theorem.
@@
Refer to Section S8.3 of the Supplementary Material for a proof of the theorem.
@@
Consistency of these estimators under weak stability is established in Section S8.3 of the Supplementary Material (as a part of proof of Theorem 3). Substitution of estimators (21) into (17) yields the following conservative estimator of $\mathrm{V}^{\mathrm{IPW}}$ under weak stability when $p_{1}^{\star},p_{2}^{\star}$ and $\widetilde{p}$ are unknown:
$\widetilde{V^{\text{IPW}}_{\text{weak}}}=\widetilde{m}_{0}^{2}\left(\frac{1}{N}\sum_{i=1}^{N}\frac{1}{1-p_{i}}-1\right)+\widetilde{m}_{1}^{2}\left(\frac{1}{N}\sum_{i=1}^{N}\frac{1}{p_{i}}-1\right)+2\widetilde{m}_{0}\widetilde{m}_{1},$ (22)
@@
Since we have constructed conservative variance estimators of $\mathrm{V}^{\mathrm{IPW}}$, for any target level $\alpha\in(0,1)$, an asymptotically conservative confidence interval for $\bar{\tau}$, that is, one with coverage at least $(1-\alpha)$ can be constructed as
$\lim_{N\rightarrow\infty}\mathbb{P}\left(\bar{\tau}\in\left[\widehat{\tau}_{\textsc{IPW}}-z_{1-\alpha/2}\sqrt{\frac{\widehat{V}}{N}},\;\;\widehat{\tau}_{\textsc{IPW}}+z_{1-\alpha/2}\sqrt{\frac{\widehat{V}}{N}}\right]\right)\geq 1-\alpha,$
where $\widehat{V}=\widehat{V^{\text{IPW}}_{\text{strong}}}$ or $\widehat{V^{\text{IPW}}_{\text{weak}}}$, and $z_{1-\alpha/2}$ denotes the $(1-\alpha/2)$th quantile of the standard normal distribution. Moreover, if the potential outcomes satisfy the log-additive treatment-effect (4), the inequality holds with equality, yielding asymptotically exact coverage.
@@ 4.2 The AIPW-type estimator
The proof of Theorem 4 requires a different strategy from that of Theorem 1, since the AIPW estimator is not directly amenable to martingale central limit theorem. To handle this, we introduce a proxy estimator, which is interesting in its own right, that is analytically more tractable and can be expressed as sum of a martingale difference sequence, allowing the martingale central limit theorem to establish its asymptotic normality. The key step is then to show that the difference between the proxy and the actual AIPW estimator is asymptotically negligible, using variance bounds and Hájek’s lemma (see Lemma 6 of the Supplementary Material). This ensures that the asymptotic distribution of the AIPW estimator coincides with that of the proxy, yielding the stated central limit theorem with variance given in Theorem 4. Refer to Section S8.4 of the Supplementary Material for detailed proof of the above theorem.
@@
Note that $\mathrm{V}^{\mathrm{AIPW}}\leq\mathrm{V}^{\mathrm{IPW}}$; that is, $\widehat{\tau}_{\textsc{AIPW}}$ is more efficient than $\widehat{\tau}_{\textsc{IPW}}$. This fact clearly establishes the superiority of the AIPW estimator over the IPW estimator in finite-population design-based inference under the adaptive assignment mechanism defined in (7). However, the corresponding variance estimators (defined later for AIPW estimators) may not inherit this ordering. See Section S10 of the Supplementary Material for a detailed discussion.
@@
A proof of this theorem is provided in Section S8.5 of the Supplementary Material. A similar remark as in Remark 5 applies.
@@
As discussed previously, the cross-moment term $\sigma_{01}$ cannot be estimated without additional assumptions, since it depends on both potential outcomes for any unit. We therefore bound it from above using the Cauchy–Schwarz inequality. Consequently, asymptotic variance $\mathrm{V}^{\text{AIPW}}_{\text{weak}}$ can be conservatively estimated by:
@@
As in the case of strong stability, this estimator is consistent when the potential outcomes satisfy generalized treatment effect additivity as defined in Definition 1. The above discussion is summarized in the following theorem.
@@
Refer to Section S8.6 of the Supplementary Material for a proof of the theorem.
@@
As in Section 4.1, for any target level $\alpha\in(0,1)$, asymptotically conservative confidence intervals for $\bar{\tau}$ may be constructed for $\widehat{\tau}_{\textsc{AIPW}}$ using the variance estimators (26) and (30) for strongly stable and weakly stable designs, respectively. If the limiting values of the probabilities are unknown, their counterparts suggested in Remark 9 may be used. All of these intervals are asymptotically conservative, but attain exact asymptotic coverage when the potential outcomes satisfy generalized treatment effect additivity (3).
@@ 5 Some illustrative applications
In this section, we illustrate Theorems 1-6 through two popular adaptive designs: a strongly stable design, Wei’s adaptive coin design [Wei1978-gd], and a weakly stable design, Efron’s biased coin design [Efron1971-yh]. We further complement the theoretical results with numerical simulations that demonstrate the validity of our approach.
@@ 5.1 Wei’s Adaptive Coin Design
We begin with Wei’s adaptive coin design [Wei1978-gd], which reduces relative imbalance between treatment and control allocations. Formally, let $m_{k}$ and $n_{k}$ denote, respectively, the numbers of treatment and control units among the first $k$ subjects. Define the treatment-control imbalance up to the $k$th assignment as $D_{k}=m_{k}-n_{k}=2m_{k}-k$, and the corresponding normalized imbalance $R_{k}=D_{k}/k,$ which measures the average difference between the treatment and control groups up to stage $k$. The $i$th subject is assigned to treatment with probability
@@
Intuitively, when the trial is in its early stages, the number of units in each group can differ substantially in relative terms; the design then shifts $p_{i}$ away from $\tfrac{1}{2}$ to favor the smaller group and reduce imbalance. As the sample size grows, any absolute difference in group sizes becomes small relative to the total number of units, causing $R_{i-1}$ to shrink and $p_{i}$ to converge to $\tfrac{1}{2}$. The following Lemma, which is a direct consequence of [Wei1978-gd, Theorem 1], establishes strong stability of the truncated version of Wei’s design (33), making Theorems 1245 directly applicable.
@@
See Section S9.1 of the Supplementary Material for the proof of Lemma 1.
Substituting $p^{\star}=\tfrac{1}{2}$ into the expressions for $\mathrm{V}^{\text{IPW}}_{\text{strong}}$ and $\mathrm{V}^{\text{AIPW}}_{\text{strong}}$ in Theorems 1 and 4 yields the limiting variances of the IPW and AIPW estimators, $\widehat{\tau}_{\textsc{IPW}}$ and $\widehat{\tau}_{\textsc{AIPW}}$, respectively, under Wei’s design:
@@
Since $p^{\star}$ is known, it can be directly plugged into the variance estimators $\widehat{V^{\text{IPW}}_{\text{strong}}}$ in (14) and $\widehat{V^{\text{AIPW}}_{\text{strong}}}$ in (26), yielding the following conservative estimators for the IPW and AIPW variances:
@@
where $\widehat{m}^{2}_{\ell}$ and $\widehat{\sigma}^{2}_{\ell}$, $\ell\in\{0,1\}$, are as defined in (15) and (27), respectively. We can now use these variance estimators to construct conservative confidence intervals for $\bar{\tau}$. Recall that the interval based on $\widehat{V^{\text{IPW}}_{\text{Wei}}}$ attains exact asymptotic coverage when the potential outcomes satisfy additivity on the log scale (4), whereas the interval based on $\widehat{V^{\text{AIPW}}_{\text{Wei}}}$ attains exact asymptotic coverage under generalized treatment effect additivity as defined in (3).
Next, we evaluate the performances of the IPW and AIPW estimators under Wei’s adaptive coin design through simulation studies. We consider three types of potential outcome matrices: (a) A general, non-additive setting, in which the potential outcomes $(Y_{i}(0),Y_{i}(1))$ are drawn from a bivariate normal distribution with mean vector $(0,1)^{\text{T}}$, variance vector $(1,1)^{\text{T}}$ and correlation 0.3, with support restricted to [3,3][-3,3] to ensure bounded outcomes, (b) an additive setting satisfying (2), where the control potential outcomes are drawn from a normal distribution with mean 0 and variance 1, truncated to [3,3][-3,3], and the treatment outcomes are defined by $Y_{i}(1)=Y_{i}(0)+\tau$ with $\tau=10$; and (c) a log-additive setting satisfying (4), in which the control potential outcomes are drawn from a normal distribution with mean 10 and variance 1, truncated to $[7,13]$, and the treatment outcomes are defined by $Y_{i}(1)=c\,Y_{i}(0)$ with $c=2$. Potential outcomes generated once are held fixed.
Treatment assignments $\mathbf{K}$ are generated according to Wei’s sequential randomization scheme, with assignment probabilities $p_{i}=f(R_{i-1})=\left(1-R_{i-1}\right)/2=1-m_{i-1}/(i-1)$, where $m_{i-1}$ denotes the number of units assigned to the treatment among the first (i1)(i-1) subjects, and the truncation parameter is set to $\delta=0.01$. Each simulation involves $N=5000$ units and is repeated $M=20000$ times. For each replication, confidence intervals are constructed as proposed, and empirical coverage is evaluated across 20 nominal levels ranging from 0.75 to 0.99.
(a) IPW estimator: Non-Additive
(b) IPW estimator: Additive
(c) IPW estimator: Log-Additive
(d) AIPW estimator: Non-Additive
(e) AIPW estimator: Additive
(f) AIPW estimator: Log-Additive
@@
Table 1 reports additional diagnostic metrics across sample sizes. Specifically, we report the Monte Carlo bias, computed as the difference between the average estimator across $M=2000$ simulation replications and the true treatment effect, and the variance ratio, defined as the ratio of the empirical Monte Carlo variance to the average estimated variance across replications. The variance ratio evaluates the calibration of the proposed variance estimators, with values close to one indicating accurate variance estimation.
While the coverage and confidence interval length plots are based on $M=20000$ Monte Carlo replications to ensure stable estimation of tail probabilities, the bias and variance ratio diagnostics in Table 1 are computed using $M=2000$ replications, which are sufficient for accurate estimation of first- and second-moment quantities. While coverage and confidence interval length results for $N=5000$ are presented graphically, Table 1 summarizes Monte Carlo bias and variance ratio diagnostics for $N=500$ and $N=10000$.
Several patterns emerge. First, the Monte Carlo bias is negligible across sample sizes, confirming the unbiasedness properties established theoretically. Second, the variance ratios exhibit mild finite sample deviations at $N=500$, particularly under the log-additive design, but generally move closer to one as $N$ increases to $10000$. This behavior is consistent with the theoretical results and illustrates convergence towards the theoretical large sample regime.
Table 1: Finite sample bias and variance ratio for $N=500$ and $N=10000$ under Wei’s design.
Bias Variance Ratio
Model Estimator N=500 N=10000 N=500 N=10000
Additive IPW 0.013 85-0.013\,85 0.002 110.002\,11 $0.9317$ $0.8600$
AIPW 0.005 9220.005\,922 0.000 549 30.000\,549\,3 $1.1050$ $0.9046$
Non-additive IPW 0.001 2380.001\,238 0.001 2040.001\,204 $0.9379$ $0.8553$
AIPW 0.002 9930.002\,993 0.001 0140.001\,014 $1.2190$ $0.9885$
Log-additive IPW 0.040 29-0.040\,29 0.005 6650.005\,665 $1.0670$ $1.0140$
AIPW 0.005 5260.005\,526 0.000 559 10.000\,559\,1 $0.9561$ $1.0090$